# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2

### v3

@@ 4 Experiments
In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks. All experiments are conducted on NVIDIA A100-80G GPUs. To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages. This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization. We select Chronos as the TS embedder since it has not been pre-trained on HAR-specific data, allowing us to evaluate the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
@@ 4.1 Datasets
To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024). These datasets vary widely in subject counts, sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments. SensorQA (see Appendix A.7) is built on these five benchmarks and supports both alignment and activity classification. All datasets are publicly available, containing no personally identifiable information, thus posing minimal ethical or privacy concerns.
@@ Results.
Table 1 compares SensorLLM and GPT-4o on the Sensor Data Understanding task. SensorLLM consistently outperforms GPT-4o across all metrics, with outputs that align more closely with ground truth and exhibit stronger trend understanding and coherence. In contrast, GPT-4o often struggles with complex numerical data and trend detection (Yehudai et al., 2024). When serving as an evaluator, GPT-4o also indicates a preference for SensorLLM. Human evaluation, conducted on shorter sequences, likewise favors SensorLLM. Performance on CAPTURE-24 is comparatively lower, likely due to longer sequences being trained under fixed parameters. Overall, these results validate the effectiveness of our alignment method in enabling LLMs to interpret complex TS data. Qualitative examples are provided in Appendix A.10.
@@ 4.3 Human Activity Recognition
We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023a), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021). We also include Chronos+MLP and GPT4TS (Zhou et al., 2023a) for a more comprehensive comparison. Full baseline details are in Appendix A.8.
@@ Results.
Table 2 reports F1-macro and accuracy scores (%) averaged over five runs. SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, slightly behind Attend. It shows notable gains on challenging datasets such as CAPTURE-24 and MHealth, demonstrating strong performance in real-world and long-sequence settings. Compared to Chronos+MLP, which uses the same TS encoder, SensorLLM significantly improves both F1-macro and accuracy, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data.
Strong results on both F1-macro and accuracy indicate that SensorLLM performs well in both overall prediction and per-class balance, showing robust generalization across diverse sensor configurations, activity types, and data collection environments.
@@ 5 Ablation Studies
To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM ’s performance with and without prompts. As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture. This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns. The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy. More broadly, the ability to jointly process sensor data and textual prompts underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
@@
To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b—a lighter variant built with Chronos-base and LLaMA3.2-3b. Experiments were conducted on USC-HAD, UCI-HAR, and MHealth. As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b, reflecting the trade-off between model size and accuracy. Nevertheless, it remains competitive—outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR. These results suggest that SensorLLM-3b provides a strong balance between efficiency and performance, making it a viable choice for real-world, resource-limited applications.
@@ 6 Conclusions
We present SensorLLM, a multimodal framework that aligns sensor data with natural language through a QA format at a human-perception level, moving beyond machine-level alignment. It effectively captures complex sensor patterns, achieves strong performance on HAR tasks, and generalizes well without requiring dataset-specific alignment. Experiments demonstrate its robustness across variable-length sequences, multivariate inputs, and textual metadata. To support future research, we release our code and the SensorQA dataset, constructed from five public HAR benchmarks, to advance time-series and language integration, particularly in low-resource domains.
@@ 7 Limitations
Classifier-Based Design.
To ensure fair comparisons with existing HAR models, we adopt a classifier for downstream tasks rather than fully leveraging the LLM’s generative capabilities. While our results demonstrate that the Sensor–Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may limit adaptability to new activity categories. Although zero-shot adaptability is a valuable direction, we did not explore it here due to the lack of comparable baselines. To the best of our knowledge, no prior work supports generalization to unseen activity classes under variable-length and variable-channel sensor input, as our framework does, making fair comparison on zero-shot settings infeasible at this stage. Future work could explore generative or prompt-based approaches to support broader applications such as activity discovery or open-set recognition.
@@
Our alignment focuses on mapping sensor data to trend-descriptive text, demonstrating clear benefits for LLM-based HAR. However, human-intuitive descriptions of sensor data extend beyond trend changes—incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data. Future research could investigate whether aligning text with alternative sensor characteristics improves time-series reasoning. This could expand the potential of multimodal NLP applications in sensor-driven tasks beyond activity recognition.
@@ A.1 More related work
LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting. Xue and Salim (2023) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings. Gruver et al. (2023) explores encoding time-series as numerical strings, allowing LLMs like GPT-3 and LLaMA-2 to perform zero-shot forecasting, matching or surpassing the performance of specialized models, while highlighting challenges in uncertainty calibration due to model modifications like RLHF. Zhou et al. (2023b) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks. Jin et al. (2024b) highlights the transformative potential of LLMs for time-series analysis by integrating language models with traditional analytical methods. Jin et al. (2024a) introduces a reprogramming framework that aligns time-series data with natural language processing capabilities, enabling LLMs to perform time-series forecasting without altering the core model structure. Cao et al. (2024) presents TEMPO, a generative transformer framework based on prompt tuning, which adapts pre-trained models for time-series forecasting by decomposing trends, seasonality, and residual information. Sun et al. (2024b) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications. Chang et al. (2024) develops a framework that enhances pre-trained LLMs for multivariate time-series forecasting through a two-stage fine-tuning process and a novel multi-scale temporal aggregation method, outperforming traditional models in both full-shot and few-shot scenarios. Finally, Liu et al. (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
LLMs for Human Activity Recognition.
While LLMs like ChatGPT have demonstrated remarkable performance in various NLP tasks, their effectiveness in HAR remains limited due to challenges in interpreting sensor data. These models often struggle to distinguish between activities that share similar objects, requiring more advanced prompt engineering to highlight activity-specific details. Xia et al. (2023) proposed an unsupervised approach to HAR using ChatGPT, leveraging two-stage prompts to infer activities from object sequences without manual descriptions. The method demonstrates superior performance on three benchmark datasets, marking a significant advancement in applying language models to activity recognition tasks. Similarly, Ji et al. (2024) explored LLMs for zero-shot HAR using raw IMU data, showing that GPT-4 can outperform both traditional and deep learning models in simple HAR tasks without domain-specific adaptations, highlighting LLMs’ potential in sensor-based systems.
A.2 SensorQA Generation
@@
For SensorQA, we generate text data from sensor readings using predefined sentence templates (Tables 6, 7, 8). These templates are randomly selected to create diverse question-answer (QA) pairs. To enhance variability, we employ GPT-4o to generate synonymous variations. Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
@@ A.6 Datasets
A.7 SensorQA
We introduce SensorQA, a novel question-answering dataset designed to align time-series sensor data with human-interpretable natural language. Each sample pairs a segment of sensor input with natural language questions and answers, capturing trends, patterns, or activity-level semantics.
Sensor-Language Alignment Stage
of SensorQA focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
Task-Aware Tuning Stage
focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage. This stage of SensorQA contains statistical information from each sensor channel as part of the input representation.
Dataset Stage 1 Stage 2
Train Test Train Test
USC-HAD 300,744 58,704 22,790 4,555
UCI-HAR 128,292 25,932 7,352 2,947
PAMAP2 738,666 271,674 14,163 5,210
MHealth 283,020 60,780 4771 2,039
CAPTURE-24 72,714 35,688 61,327 30,138
Table 12: Training and testing sample counts for Stage 1 and Stage 2 across datasets of SensorQA.
A.8 Baselines for Task-Aware Tuning Stage
@@
A.9 Evaluation Metrics for Task-Aware Tuning Stage
@@
[…]
@@
A.10 Sensor-Language Alignment Stage Output Examples

### v4

@@ 4 Experiments
In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks. All experiments are conducted on NVIDIA A100-80G GPUs. To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages. This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization. We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
@@ 4.1 Datasets
To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024). These datasets vary widely in subject counts, sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments. All datasets are publicly available, containing no personally identifiable information, thus posing minimal ethical or privacy concerns.
@@ Results.
Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics. BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps. SBERT and SimCSE can capture factual correctness or deeper semantic similarities. Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth. GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024). Human evaluation also favors SensorLLM, particularly for shorter sequences. CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters. Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence. Appendix A.9 provides qualitative examples of outputs from both models.
Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
[…]
@@ 4.3 Human Activity Recognition
We benchmark SensorLLM against 11 baselines across two categories: (i) TS models—Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021), NS-Transformer (Liu et al., 2022), PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023), and iTransformer (Liu et al., 2024c); (ii) HAR models—DeepConvLSTM (Ordóñez and Roggen, 2016), DeepConvLSTMAttn (Murahari and Plötz, 2018), and Attend (Abedin et al., 2021). We also include Chronos+MLP and GPT4TS (Zhou et al., 2023a) for a more comprehensive comparison. Full baseline details are in Appendix A.7.
@@ Results.
Table 2 reports the F1-macro and accuracy scores (%) averaged over five random runs. SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, indicating strong performance across diverse sensor settings. On CAPTURE-24, SensorLLM reaches 48.6% F1-macro, a +5.0% absolute gain over Attend (43.6%). It also leads on USC-HAD (61.2%, +1.0%), PAMAP2 (86.2%, +1.6%), and MHealth (89.4%, +6.0%), establishing state-of-the-art performance on these datasets. On UCI-HAR, SensorLLM achieves 91.2%, slightly below Attend (93.2%).
However, Chronos+MLP offers only a marginal improvement over iTransformer on UCI-HAR (82.2% vs. 81.8%), suggesting that Chronos embeddings alone are of limited utility for this dataset. In contrast, using the same time-series encoder (Chronos), SensorLLM substantially boosts both F1-macro and accuracy, with especially large gains on challenging, real-world, long-sequence settings such as CAPTURE-24 and MHealth, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data. Overall, the consistent gains in both F1-macro and accuracy indicate balanced per-class performance alongside strong overall prediction, with robust generalization across sensor configurations, activity types, and data-collection environments.
@@ 5 Ablation Studies
To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM’s performance with and without prompts. As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture. This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns. The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy. More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
@@
To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b, a lighter variant built with Chronos-base and LLaMA3.2-3b. Experiments were conducted on USC-HAD, UCI-HAR, and MHealth. As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b, reflecting the trade-off between model size and accuracy. Nevertheless, it remains competitive, outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR. These results suggest that SensorLLM-3b provides a strong balance between efficiency and performance, making it a viable choice for real-world, resource-limited applications.
@@ 6 Conclusions
We present SensorLLM, the first multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment. SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks. Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata. Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment. This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data reasoning. We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
@@ 7 Limitations
Choice of Classifier.
It is worth noting that the primary contribution of our work lies in proposing a novel TS-text alignment strategy and rigorously validating its effectiveness. To ensure fair comparisons with existing HAR baselines, we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities. While this design allows rigorous validation of our proposed alignment method, it also imposes limitations: relying on a fixed-class classifier may constrain adaptability to new activity categories and does not fully leverage the reasoning potential of LLMs. Although our framework is compatible with other downstream heads, such as a language modeling (LM) head for prompt-based generation, we leave this exploration to future work. Investigating generative or zero-shot approaches could enable broader applications, including activity discovery and open-set recognition.
@@
Our alignment focuses on mapping sensor data to trend-descriptive text, demonstrating clear benefits for LLM-based HAR. However, human-intuitive descriptions of sensor data extend beyond trend changes, incorporating frequency-domain features, periodicity, and higher-order patterns may further enhance an LLM’s ability to interpret time-series data. Future research could investigate whether aligning text with alternative sensor characteristics improves time-series reasoning. This could expand the potential of multimodal LLM applications in sensor-driven tasks beyond activity recognition.
@@ A.1 More related work
LLMs have achieved remarkable success in text-related tasks, and their utility has expanded into time-series forecasting. Xue and Salim (2024) presents PromptCast, which redefines time-series forecasting as a natural language generation task by transforming numerical inputs into textual prompts, enabling pre-trained language models to handle forecasting tasks with superior generalization in zero-shot settings. Gruver et al. (2023) explores encoding time-series as numerical strings, allowing LLMs like GPT-3 and LLaMA-2 to perform zero-shot forecasting, matching or surpassing the performance of specialized models, while highlighting challenges in uncertainty calibration due to model modifications like RLHF. Zhou et al. (2023a) demonstrates that pre-trained language and image models, such as a Frozen Pretrained Transformer (FPT), can be adapted for diverse time-series tasks like classification, forecasting, and anomaly detection, leveraging self-attention mechanisms to bridge the gap between different data types and achieving state-of-the-art performance across various tasks. Jin et al. (2024b) highlights the transformative potential of LLMs for time-series analysis by integrating language models with traditional analytical methods. Jin et al. (2024a) introduces a reprogramming framework that aligns time-series data with natural language processing capabilities, enabling LLMs to perform time-series forecasting without altering the core model structure. Cao et al. (2024) presents TEMPO, a generative transformer framework based on prompt tuning, which adapts pre-trained models for time-series forecasting by decomposing trends, seasonality, and residual information. Sun et al. (2024) proposes TEST, an innovative embedding technique that integrates time-series data with LLMs through instance-wise, feature-wise, and text-prototype-aligned contrast, yielding improved or comparable results across various applications. Chang et al. (2024) develops a framework that enhances pre-trained LLMs for multivariate time-series forecasting through a two-stage fine-tuning process and a novel multi-scale temporal aggregation method, outperforming traditional models in both full-shot and few-shot scenarios. Finally, Liu et al. (2024b) introduces UniTime, a unified model that leverages language instructions and a Language-TS Transformer to handle multivariate time series across different domains, demonstrating enhanced forecasting performance and zero-shot transferability.
A.2 Sensor–Text Data Pair Generation
@@
We generate text data from sensor readings using predefined sentence templates (Tables 6, 8, 9). These templates are randomly selected to create diverse question-answer (QA) pairs. To enhance variability, we employ GPT-4o to generate synonymous variations. Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
@@
Dataset Stage 1 Stage 2
Train Test Train Test
USC-HAD 300,744 58,704 22,790 4,555
UCI-HAR 128,292 25,932 7,352 2,947
PAMAP2 738,666 271,674 14,163 5,210
MHealth 283,020 60,780 4771 2,039
CAPTURE-24 72,714 35,688 61,327 30,138
Table 7: Training and testing sensor-text QA pairs counts for Stage 1 and Stage 2.
@@
Sensor-Language Alignment Stage
focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
Task-Aware Tuning Stage
focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage. This stage contains statistical information from each sensor channel as part of the input representation.
@@ A.6 Datasets
A.7 Baselines for Task-Aware Tuning Stage
@@
A.8 Evaluation Metrics for Task-Aware Tuning Stage
@@
$\text{F1-macro}=\frac{1}{C}\sum_{i=1}^{C}\text{F1}_{i}$ (8)
where $C$ is the total number of classes, and $\text{F1}_{i}$ is the F1 score for class $i$. The F1 score for each class is calculated as:
$\text{F1}_{i}=\frac{2\times\text{Precision}_{i}\times\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (9)
@@
A.9 Sensor-Language Alignment Stage Output Examples